If np6=3np5 find value of n
Answer by ikleyn(52778)

.
There is nothing easier.

nP6 = n*(n-1)*(n-2)*(n-3)*(n-4)*(n-5)   (6 factors).


nP5 = n*(n-1)*(n-2)*(n-3)*(n-4)         (5 factors)


The condition gives you this equation


     n*(n-1)*(n-2)*(n-3)*(n-4)*(n-5) = 3 * n*(n-1)*(n-2)*(n-3)*(n-4).



Cancel   n*(n-1)*(n-2)*(n-3)*(n-4)   in both sides.


You will get


     n-5 = 3.


Answer.  n = 8.